OpenStudy (anonymous):

what is the secant of theta= 5pi/4?

OpenStudy (dumbcow):

\[\sec(\theta)=\frac{5\pi}{4}\] \[\frac{1}{\cos(\theta)} = \frac{5\pi}{4}\] \[1 = \frac{5\pi}{4}\cos(\theta)\] \[\cos(\theta) = \frac{4}{5 \pi}\] \[\theta = \cos^{-1}(\frac{4}{5\pi}) \approx 75.25^{o}\]